OpenStudy (perii224):

How do you solve this? Solve for x. y = + 8

OpenStudy (anonymous):

I don't think that's the full equation...

OpenStudy (perii224):

Its y= x/3 + 8

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

Subtract 8 from both sides.

OpenStudy (skullpatrol):

Multiply both sides by 3 and you're finished :)

OpenStudy (skullpatrol):

The goal is to isolate x on one side of the equation.

OpenStudy (perii224):

Is this the correct answer? x = 3(y – 8)

OpenStudy (perii224):

x = 3(y-8)

OpenStudy (skullpatrol):

3(y-8) = x

OpenStudy (perii224):

Ok Got It! Thanks :)

OpenStudy (skullpatrol):

Yes, that is correct :)

OpenStudy (skullpatrol):

Thanks for trying to learn.